# Contextual Status Updating: How to Maximize Situational Awareness?

### v1

@@ Abstract
In this study, we consider a problem of remote safety monitoring, where a monitor pulls status updates from multiple sensors monitoring several safety-critical situations. Based on the received updates, multiple estimators determine the current safety-critical situations. Due to transmission errors and limited channel resources, the received status updates may not be fresh, resulting in the possibility of misunderstanding the current safety situation. In particular, if a dangerous situation is misinterpreted as safe, the safety risk is high. We study the joint design of transmission scheduling and estimation for multi-sensor, multi-channel remote safety monitoring, aiming to minimize the loss due to the unawareness of potential danger. We show that the joint design of transmission scheduling and estimation can be reduced to a sequential optimization of estimation and scheduling. The scheduling problem can be formulated as a Restless Multi-armed Bandit (RMAB) , for which it is difficult to establish indexability. We propose a low-complexity Maximum Gain First (MGF) policy and prove it is asymptotically optimal as the numbers of sources and channels scale up proportionally, without requiring the indexability condition. We also provide an information-theoretic interpretation of the transmission scheduling problem. Numerical results show that our estimation and scheduling policies achieves higher performance gain over periodic updating, randomized policy, and Maximum Age First (MAF) policy.
@@ I Introduction
One challenge to efficiently utilize the state information in real-time is the limited capacity of the communication medium. Conversely, depending on the surrounding situation, some processes can be more significant to monitor than others For example, in autonomous driving, a self-driving car deciding to change lanes need to prioritize real-time information about vehicles in the adjacent lane over those in its current lane. To address this issue, it is important to quantify the significance, or the value, of status updates, for safety monitoring.
In this paper, we answer this question for a pull-based status updating system where multiple sensors monitor the status of several safety-critical situations. A central scheduler requests updates from sensors. Due to transmission errors and limited channel resources,, the received updates may not be fresh. One performance metric that characterizes data freshness is the Age of Information (AoI) [5]. Let $U(t)$ be the generation time of the freshest received observation by time $t$. AoI, as a function of $t$, is defined as $\Delta(t)=t-U(t)$ which exhibits a linear growth with time $t$ and drops down to a smaller value whenever a fresher observation is delivered. However, AoI only captures the timeliness of the information, but not its significance. Hence, relying solely on AoI-based decision making is not sufficient, particularly, in safety-critical scenarios where misunderstanding about the situation can lead to significant performance loss. To address this, we employ multiple estimators in the receiver that estimate the safety situations, and by incorporating a loss function, we measure the significance of the estimated status update for safety monitoring. One important property of the loss function is: If a dangerous situation is incorrectly estimated as safe, the loss will be significantly higher because of the high risk of damage. Conversely, if the safe situation is incorrectly estimated as dangerous, the loss will be small. This is because even if the estimate is incorrect, the risk of damage is small.
@@
  • Numerical results illustrate that our scheduling policy achieves significant performance gain compared to periodic updating policy, randomized policy, and Maximum Age First policy (see Section VII). Because our policy utilizes the knowledge of the latest received signal, it illustrates good performance compared to the other policies that ignore this knowledge.
@@ II Related Work
AoI in Sampling and Scheduling: There exist numerous papers on AoI-based sampling and scheduling [22, 16, 20, 19, 41, 42, 14, 28, 43, 31, 32, 44]. Authors in [43] studied an AoI minimization problem under a pulling model that considers replicated requests to the server. In [22], sampling policies for optimizing non-linear AoI functions were studied. AoI minimization for single-hop networks was studied in [45]. A joint sampling and scheduling problem to minimize monotonic AoI functions were considered in [30]. A Whittle index-based scheduling algorithm to minimize AoI for stochastic arrivals was considered in [46]. Authors in [18] proposed a Whittle index policy for minimizing non-decreasing AoI functions. In [16], the authors proposed a Whittle index-based scheduling policy to minimize the UoI modeled as Shanon entropy. Optimal scheduling policies for both single and multi-source systems were studied and a Whittle index policy was proposed for multi-source case in [20]. The optimal sampling policies for Gauss-Markov processes were studied in [31, 32, 44] where the estimation error becomes a monotonic function of age in signal-agnostic scenarios and the associated problem for minimizing age-penalty functions were reported. A Whittle index policy for continuous-time Gauss Markov processes for both signal-aware and signal-agnostic scenarios was reported in [19]. Besides Whittle index-based policies that require an indexability condition, non-indexable scheduling policies were also studied in [41, 42, 14, 15, 28]. In this paper, because of the complicated nature of state transition along with erasure channels, indexability is very difficult to establish. However, we provide a “Maximum Gain First Policy” developed in [28, 15]. The comparison of our model with [28, 15] is that our model considers erasure channels, signal observation, and general loss functions. Our scheduling policy is designed for a pull-based communication model where the scheduling decisions are based on the latest received observation and its AoI. We further proved that the developed policy is asymptotically optimal.
@@ III-A System Model
Due to channel sharing among the agents and transmission errors, the received information may not always be up to date. The most recent packet available at time $t$ is denoted as $X_{n,t-\Delta_{n}(t)}$, which was originally generated at time $t-\Delta_{n}(t)$. In remote estimation systems, the time difference $\Delta_{n}(t)$ between the packet’s generation time and the current time $t$, $t-\Delta_{n}(t)$ is known as age of information (AoI) [5, 48, 32, 31]. AoI quantifies the freshness of status updates received from agent $n$ and evolves according to:
$\Delta_{n}(t+1)=\begin{cases}&1,\quad\text{if }\mu_{n}(t)=1\text{and }\gamma_{n}(t)=1,\\ &\Delta_{n}(t)+1,\text{otherwise.}\end{cases}$ (2)
@@ III-B Safety Level Estimator
The loss values and the number of safety levels can be flexibly designed to fit different applications. By this, the loss function $L_{n}(y,\hat{y})$ can be adjusted to meet specific situational awareness requirements, ensuring adaptability across various contexts. A key advantage of the general loss function $L_{n}(\cdot,\cdot)$ is its ability to address safety risks arising from unawareness of potential danger—something traditional loss functions, such as 0-1 loss, quadratic loss, and logarithmic loss [report, Appendix A], fail to capture.
@@
$\mathbf{H}_{n,t}=\big{(}X_{n,\tau-\Delta_{n}(\tau)},\Delta_{n}(\tau),\mu_{n}(\tau-1),\gamma_{n}(\tau-1)\big{)}_{\tau=0}^{t},$ (3)
@@ IV Problem Simplification
$H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})=$
$\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\delta}}(\delta,x)\times$
$\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]},$ (13)
@@ V Scheduling Policy Design for Solving (14)-(15)
Problem (14)-(15) is a Restless Multi-armed Bandit (RMAB). The problem (14)-(15) is called “restless” because the state and the penalty of each bandit $n$ continues to evolve over time, regardless of whether the agent $n$ is selected for transmission [51]. Because we are able to simplify the original problem (4)-(5), the RMAB (14)-(15) has a reduced state space compared to (4)-(5). However, even with the reduced state space, solving RMAB problems and finding optimal solutions are significantly challenging. solving RMAB (14)-(15) remains a significant challenge. A Whittle index policy is known to be an efficient approach to solving RMAB problems which requires to satisfy a condition called indexability[51], [52]. A key challenge in solving problem (14)-(15) is that indexability is very difficult to establish. This difficulty arises due to the following reasons: (i) The state of each bandit of RMAB (14)-(15) exhibits a complicated transition, (ii) the transmission channels are unreliable, and (iii) the expected penalty associated with each bandit can be non-monotonic function of the AoI while most of the previous studies considered monotonic penalty functions of AoI [53, 22, 54, 18]. Hence, (14)-(15) is a more challenging problem than the problems studied in [53, 22, 54, 18, 16, 15]. However, we are able to develop a Maximum Gain First (MGF) policy that does not need to satisfy indexability.
@@ V-A Relaxation and Lagrangian Decomposition
$\bar{L}(\lambda)=$ $\!\inf_{\pi\in\Pi}\!\limsup_{T\to\infty}\!\!\bigg{[}\sum_{n=1}^{N}\!\frac{1}{T}\!\!\sum_{t=0}^{T-1}\!H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda\bigg{(}\mathbb{E}\bigg{[}\sum_{n=0}^{N}\mu_{n}(t)\bigg{]}-M\bigg{)}\bigg{]}.$ (25)
The term $\frac{1}{T}\sum_{t=0}^{T-1}\sum_{n=0}^{N}\lambda M$ in (V-A) does not depend on policy $\pi$ and hence can be removed. For a given $\lambda$, problem (V-A) can be decomposed into $(N+1)$ separated sub-problems and each sub-problem associated with agent $n$ is formulated as
$\bar{L}_{n}(\lambda)=$ $\inf_{\pi_{n}\in\Pi_{n}}\!\!\limsup_{T\to\infty}\bigg{[}\frac{1}{T}\sum_{t=0}^{T-1}H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda\mathbb{E}[\mu_{n}(t)]\bigg{]},$ (26)
where $\bar{L}_{n}(\lambda)$ is the optimum value of (V-A), $\pi_{n}=(\mu_{n}(0),\mu_{n}(1),\ldots)$ is the sub-scheduling policy for agent $n$, and $\Pi_{n}$ is the set of all causal sub-scheduling policies of agent $n$. Problem (V-A) is a per-bandit problem associated with bandit $n$. On the other hand, the sub-problem associated with the dummy bandits is given by
@@ V-B Maximum Gain First (MGF) Policy
For a given transmission cost $\lambda$, the per-bandit problem (V-A) can be cast as an average-cost infinite horizon MDP with state $(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$. The state associated with each bandit $n$ is the latest received observation $X_{n,t-\Delta_{n}(t)}$ and its AoI $\Delta_{n}(t)$ at time slot $t$. The action is defined by $\mu_{n}(t)\in\{0,1\}$ which denotes the scheduling decision for agent $n$ at every time slot $t$. Each MDP associated with each bandit $n$ has two actions: active and passive. If a packet from agent $n$ is requested and submitted to a channel at time slot $t$, then restless bandit $n$ takes an active action at time slot $t$; otherwise, bandit $n$ is made passive at time slot $t$.
@@
We solve (V-A) by using dynamic programming [17]. The Bellman optimality equation for the MDP in (V-A) for given $\Delta_{n}(t)=\delta$ and $X_{n,t-\Delta_{n}(t)}=x$ is
@@
$Q_{n,\lambda}(\delta,x,\mu)=$
$\begin{cases}q_{n}(\delta,x)-\bar{L}_{n}(\lambda)+h_{n,\lambda}(\delta+1,x),&\!\!\!\!\!\!\!\!\!\text{if}{\thinspace}\mu=0,\\ q_{n}(\delta,x)-\bar{L}_{n}(\lambda)+(1-p_{n})h_{n,\lambda}(\delta+1,x)\\ +p_{n}\mathbb{E}[h_{n,\lambda}(1,X_{n,0})|X_{n,-\delta}=x]+\lambda,&\!\!\!\!\!\!\!\!\!\text{otherwise},\end{cases}$ (29)
@@
and (V-B) holds because $X_{n,t}$ is a time-homogeneous Markov chain. The relative value function $h_{n,\lambda}(\delta,x)$ can be computed by using relative value iteration algorithm for average-cost MDP [17].
Let $\pi_{n,\lambda}^{*}=(\mu_{n,\lambda}^{*}(1),\mu_{n,\lambda}^{*}(2),\ldots)$ be an optimal solution to (V-A). The optimal decision at time slot $t$ for bandit $n$ is given by
@@
Further discussion on a special case of single-source, single-channel (i.e., $N=M=1$) is provided in Appendix E. For this special case, we find that for single-source, single-channel, it is always better to send. Earlier studies [29, 22, 20] reported that for single-source, single-channel scenario, it is not always better to take the active action for the following two cases: (i) if the penalty function is a non-monotonic function of the age, (ii) if the transmission times are random. However, in our study, though the penalty $q_{1}(\delta,x)$ is not necessarily a monotonic function of the age, we show that it is always better to send.

### v2

@@ Abstract
In this study, we consider a problem of remote safety monitoring, where a monitor pulls status updates from multiple sensors monitoring several safety-critical situations. Based on the received updates, multiple estimators determine the current safety-critical situations. Due to transmission errors and limited channel resources, the received status updates may not be fresh, resulting in the possibility of misunderstanding the current safety situation. In particular, if a dangerous situation is misinterpreted as safe, the safety risk is high. We study the joint design of transmission scheduling and estimation for multi-sensor, multi-channel remote safety monitoring, aiming to minimize the loss due to the unawareness of potential danger. We show that the joint design of transmission scheduling and estimation can be reduced to a sequential optimization of estimation and scheduling. The scheduling problem can be formulated as a Restless Multi-armed Bandit (RMAB), for which it is difficult to establish indexability. We propose a low-complexity Maximum Gain First (MGF) policy and prove it is asymptotically optimal as the numbers of sources and channels scale up proportionally, without requiring the indexability condition. We also provide an information-theoretic interpretation of the transmission scheduling problem. Numerical results show that our estimation and scheduling policies achieves higher performance gain over random selection with queue, randomized policy, and Maximum Age First (MAF) policy.
@@ I Introduction
One challenge to efficiently utilize the state information in real-time is the limited capacity of the communication medium. Conversely, depending on the surrounding situation, some processes can be more significant to monitor than others For example, in autonomous driving, a self-driving car deciding to change lanes needs to prioritize real-time information about vehicles in the adjacent lane over those in its current lane. To address this issue, it is important to quantify the significance, or the value, of status updates, for safety monitoring.
In this paper, we answer this question for a pull-based status updating system where multiple sensors monitor the status of several safety-critical situations. A central scheduler requests updates from sensors. Due to transmission errors and limited channel resources, the received updates may not be fresh. One performance metric that characterizes data freshness is the Age of Information (AoI) [5]. Let $U(t)$ be the generation time of the freshest received observation by time $t$. AoI, as a function of $t$, is defined as $\Delta(t)=t-U(t)$ which exhibits a linear growth with time $t$ and drops down to a smaller value whenever a fresher observation is delivered. However, AoI only captures the timeliness of the information, but not its significance. Hence, relying solely on AoI-based decision making is not sufficient, particularly, in safety-critical scenarios where misunderstanding about the situation can lead to significant performance loss. To address this, we employ multiple estimators in the receiver that estimate the safety situations, and by incorporating a loss function, we measure the significance of the estimated status update for safety monitoring. One important property of the loss function is: If a dangerous situation is incorrectly estimated as safe, the loss will be significantly higher because of the high risk of damage. Conversely, if the safe situation is incorrectly estimated as dangerous, the loss will be small. This is because even if the estimate is incorrect, the risk of damage is small.
@@
  • Numerical results illustrate that our scheduling policy achieves significant performance gain compared to random selection with queue, randomized policy, and Maximum Age First policy (see Section VII). Because our policy utilizes the knowledge of the latest received signal, it illustrates good performance compared to the other policies that ignore this knowledge.
@@ II Related Work
AoI in Sampling and Scheduling: There exist numerous papers on AoI-based sampling and scheduling [22, 16, 20, 19, 41, 42, 14, 28, 43, 31, 32, 44]. Authors in [43] studied an AoI minimization problem under a pulling model that considers replicated requests to the server. In [22], sampling policies for optimizing non-linear AoI functions were studied. AoI minimization for single-hop networks was studied in [45]. A joint sampling and scheduling problem to minimize monotonic AoI functions were considered in [30]. A Whittle index-based scheduling algorithm to minimize AoI for stochastic arrivals was considered in [46]. Authors in [18] proposed a Whittle index policy for minimizing non-decreasing AoI functions. In [16], the authors proposed a Whittle index-based scheduling policy to minimize the UoI modeled as Shanon entropy. Optimal scheduling policies for both single and multi-source systems were studied and a Whittle index policy was proposed for multi-source case in [20]. The optimal sampling policies for Gauss-Markov processes were studied in [31, 32, 44] where the estimation error becomes a monotonic function of age in signal-agnostic scenarios and the associated problem for minimizing age-penalty functions were reported. A Whittle index policy for continuous-time Gauss-Markov processes for both signal-aware and signal-agnostic scenarios was reported in [19]. Besides Whittle index-based policies that require an indexability condition, non-indexable scheduling policies were also studied in [41, 42, 14, 15, 28]. In this paper, because of the complicated nature of state transition along with erasure channels, indexability is very difficult to establish. However, we provide a “Maximum Gain First Policy” developed in [28, 15]. The comparison of our model with [28, 15] is that our model considers erasure channels, signal observation, and general loss functions. Our scheduling policy is designed for a pull-based communication model where the scheduling decisions are based on the latest received observation and its AoI. We further proved that the developed policy is asymptotically optimal.
@@ III-A System Model
Due to channel sharing among the agents and transmission errors, the received information may not always be up to date. The most recent packet available at time $t$ is denoted as $X_{n,t-\Delta_{n}(t)}$, which was originally generated at time $t-\Delta_{n}(t)$. In remote estimation systems, the time difference $\Delta_{n}(t)$ between the packet’s generation time $t-\Delta_{n}(t)$ and the current time $t$ is known as age of information (AoI) [5, 48, 32, 31]. AoI quantifies the freshness of status updates received from agent $n$ and evolves according to:
$\Delta_{n}(t+1)=\begin{cases}&\!\!\!\!\!\!1,~~~~~~~~~~~\text{if}{\thinspace}\mu_{n}(t)=1~{\thinspace}\text{and}~{\thinspace}\gamma_{n}(t)=1,\\ &\!\!\!\!\!\!\Delta_{n}(t)+1,\text{otherwise.}\end{cases}$ (2)
@@ III-B Safety Level Estimator
The loss values and the number of safety levels can be flexibly designed to fit different applications. By this, the loss function $L_{n}(y,\hat{y})$ can be adjusted to meet specific situational awareness requirements, ensuring adaptability across various contexts. A key advantage of the general loss function $L_{n}(\cdot,\cdot)$ is its ability to address safety risks arising from unawareness of potential danger—something traditional loss functions, such as 0-111 loss, quadratic loss, and logarithmic loss, (See Appendix A for definitions) fail to capture.
@@
$\mathbf{H}_{n,t}=(X_{n,\tau-\Delta_{n}(\tau)},\Delta_{n}(\tau),\mu_{n}(\tau-1),\gamma_{n}(\tau-1))_{\tau=0}^{t},$ (3)
@@ IV Problem Simplification
$H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})=\sum_{(\delta,x)\in\mathbb{Z}\times\mathcal{X}_{n}}P_{\Delta_{n}(t),X_{n,t-\delta}}(\delta,x)\inf_{f_{n}(\delta,x)\in\mathcal{Y}_{n}}\mathbb{E}_{Y\sim P_{Y_{n,t}|\Delta_{n}(t)=\delta,X_{n,t-\delta}=x}}\big{[}L_{n}(Y,f_{n}(\delta,x))\big{]},$ (13)
@@ V Scheduling Policy Design for Solving (14)-(15)
Problem (14)-(15) is a Restless Multi-armed Bandit (RMAB). The problem (14)-(15) is called “restless” because the state and the penalty of each bandit $n$ continues to evolve over time, regardless of whether the agent $n$ is selected for transmission [51]. Because we are able to simplify the original problem (4)-(5), the RMAB (14)-(15) has a reduced state space compared to (4)-(5). However, even with the reduced state space, solving RMAB problems and finding optimal solutions are significantly challenging. A Whittle index policy is known to be an efficient approach to solving RMAB problems which requires to satisfy a condition called indexability[51], [52]. A key challenge in solving problem (14)-(15) is that indexability is very difficult to establish. This difficulty arises due to the following reasons: (i) The state of each bandit of RMAB (14)-(15) exhibits a complicated transition, (ii) the transmission channels are unreliable, and (iii) the expected penalty associated with each bandit can be non-monotonic function of the AoI while most of the previous studies considered monotonic penalty functions of AoI [53, 22, 54, 18]. Hence, (14)-(15) is a more challenging problem than the problems studied in [53, 22, 54, 18, 16, 15]. However, we are able to develop a Maximum Gain First (MGF) policy that does not need to satisfy indexability.
@@ V-A Relaxation and Lagrangian Decomposition
$\bar{L}(\lambda)=\!\inf_{\pi\in\Pi}\!\limsup_{T\to\infty}\bigg{[}\sum_{n=1}^{N}\!\frac{1}{T}\!\!\sum_{t=0}^{T-1}\!H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})+\lambda\bigg{(}\mathbb{E}\bigg{[}\sum_{n=0}^{N}\mu_{n}(t)\bigg{]}-M\bigg{)}\bigg{]}.$ (25)
[…]
$\bar{L}_{n}(\lambda)=\inf_{\pi_{n}\in\Pi_{n}}\!\!\limsup_{T\to\infty}\bigg{[}\frac{1}{T}\sum_{t=0}^{T-1}H_{L}(Y_{n,t}|\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})+\lambda\mathbb{E}[\mu_{n}(t)]\bigg{]},$ (26)
where $\bar{L}_{n}(\lambda)$ is the optimum value of (26), $\pi_{n}=(\mu_{n}(0),\mu_{n}(1),\ldots)$ is the sub-scheduling policy for agent $n$, and $\Pi_{n}$ is the set of all causal sub-scheduling policies of agent $n$. Problem (26) is a per-bandit problem associated with bandit $n$. On the other hand, the sub-problem associated with the dummy bandits is given by
@@ V-B Maximum Gain First (MGF) Policy
For a given transmission cost $\lambda$, the per-bandit problem (26) can be cast as an average-cost infinite horizon MDP with state $(\Delta_{n}(t),X_{n,t-\Delta_{n}(t)})$. The state associated with each bandit $n$ is the latest received observation $X_{n,t-\Delta_{n}(t)}$ and its AoI $\Delta_{n}(t)$ at time slot $t$. The action is defined by $\mu_{n}(t)\in\{0,1\}$ which denotes the scheduling decision for agent $n$ at every time slot $t$. Each MDP associated with each bandit $n$ has two actions: active and passive. If a packet from agent $n$ is requested and submitted to a channel at time slot $t$, then restless bandit $n$ takes an active action at time slot $t$; otherwise, bandit $n$ is made passive at time slot $t$.
@@
We solve (26) by using dynamic programming [17]. The Bellman optimality equation for the MDP in (26) for given $\Delta_{n}(t)=\delta$ and $X_{n,t-\Delta_{n}(t)}=x$ is
@@
$Q_{n,\lambda}(\delta,x,\mu)=\begin{cases}q_{n}(\delta,x)-\bar{L}_{n}(\lambda)+h_{n,\lambda}(\delta+1,x),&\text{if}{\thinspace}\mu=0,\\ q_{n}(\delta,x)-\bar{L}_{n}(\lambda)+(1-p_{n})h_{n,\lambda}(\delta+1,x)+p_{n}\mathbb{E}[h_{n,\lambda}(1,X_{n,0})|X_{n,-\delta}=x]+\lambda,&\text{otherwise},\end{cases}$ (29)
@@
and (29) holds because $X_{n,t}$ is a time-homogeneous Markov chain. The relative value function $h_{n,\lambda}(\delta,x)$ can be computed by using relative value iteration algorithm for average-cost MDP [17].
Let $\pi_{n,\lambda}^{*}=(\mu_{n,\lambda}^{*}(1),\mu_{n,\lambda}^{*}(2),\ldots)$ be an optimal solution to (26). The optimal decision at time slot $t$ for bandit $n$ is given by
@@
Further discussion on a special case of single-source, single-channel (i.e., $N=M=1$) is provided in Appendix F. For this special case, we find that for single-source, single-channel, it is always better to send. Earlier studies [29, 22, 20] reported that for single-source, single-channel scenario, it is not always better to take the active action for the following two cases: (i) if the penalty function is a non-monotonic function of the age, (ii) if the transmission times are random. However, in our study, though the penalty $q_{1}(\delta,x)$ is not necessarily a monotonic function of the age, we show that it is always better to send.